OpenStudy (mckenzieandjesus):

solve the system using elimination. 5x + 8y = –29 7x – 2y = –67 (–7, 9) (-10, 21/8) (–9, 2) (–1, –3)

OpenStudy (mckenzieandjesus):

@jim_thompson5910

Mehek (mehek14):

we can eliminate y

Mehek (mehek14):

multiply everything in the second equation by 4

OpenStudy (mckenzieandjesus):

7*4=28 2*4=8 right?

Mehek (mehek14):

it's actually -2*4 = -8

OpenStudy (mckenzieandjesus):

do i multiply -67 too?

OpenStudy (mckenzieandjesus):

oops yeah

OpenStudy (mckenzieandjesus):

What do i do next?

Mehek (mehek14):

yes you multiply -67 * 4

OpenStudy (mckenzieandjesus):

-268

Mehek (mehek14):

so \(\bf{5x+8y=-29\\28x-8y=-268}\)

Mehek (mehek14):

now add 5x +28x = 33x 8y + (-8y) = 0 -29 + (-268) = -279

OpenStudy (mckenzieandjesus):

ok

OpenStudy (mckenzieandjesus):

what do I do next?

Mehek (mehek14):

-29 + (-268) = -297

Mehek (mehek14):

so 33x = 297 solve for x

OpenStudy (mckenzieandjesus):

33*9=297 so x=9

Mehek (mehek14):

yes

Mehek (mehek14):

wait -297

Mehek (mehek14):

so x = -9

OpenStudy (mckenzieandjesus):

oh ok

Mehek (mehek14):

so (-9,2)

OpenStudy (mckenzieandjesus):

so its (-9, 2)

Mehek (mehek14):

yup

OpenStudy (mckenzieandjesus):

thanku!!

Mehek (mehek14):

yw :]

OpenStudy (mckenzieandjesus):

Would u know how to do this? What is the solution of the system of equations? y = –4x + 10 y = –2x – 6 (–2.67, 20.67) (–22, 8) (–2, –2) (8, –22)